exercise 9.12. tq very much

Exercise 9.12 0.05 m m=0.1 kg = 0,0001 kem los = 0.0002 kgm Determine loc (Answer: Pee = 0.0005 kgmº) b 0.03 m
Answer #1

Ans:

I​​​​​​bb​= I​​​​​​aa​​​​+ M*d​2  =0.0001+0.1*0.052 = 0.0003 kgm2. ​​I​​​​​​bb.=0.0003+0.0002= 0.0005 kgm​​​​2.   

I​​​​​​cc. =0.0005+0.01*0.032. =0.0005 kgm​2.